# 3]Sorbonne Université, CNRS, Université de Paris, Inria, Laboratoire Jacques-Louis Lions UMR 7598, 75005 Paris, France

### v1

@@ 4 Mathematical tools and techniques
In this section, we provide a concise overview of mathematical tools and techniques that have been developed and deployed to: derive PS-PIDE models of form (2) from agent-based models for the spatial spread and evolutionary dynamics of phenotype-structured cell populations (Section 4.1); analyse the qualitative properties of the solutions to such PS-PIDE models (Section 4.2); and construct numerical solutions of these models (Section 4.3).
@@
4.1 Tools and techniques to derive PS-PIDE models from agent-based models
Instead of defining a PS-PIDE model of type (2) on the basis of population-scale phenomenological assumptions, it can be desirable to first postulate an ABM that tracks the spatial and evolutionary dynamics of individual cells, and then employ limiting procedures to derive the corresponding PS-PIDE model [78]. In this way, the population-level terms that comprise the PS-PIDE can be formally linked to explicit assumptions for the cellular processes that drive the dynamics at the single-cell level.
A range of approaches have been advanced and adopted in recent decades for transitioning between ABMs and PS-PIDE models which describe simultaneously the spatial spread and evolutionary dynamics of phenotype-structured populations. These include probabilistic methods to derive PS-PIDE models as the limit of corresponding off-lattice ABMs when the number of individuals in the population tends to infinity (see e.g.  Andrade-Restrepo et al. 16, Champagnat and Méléard 65, Fontbona and Méléard 100, Leman 158), and formal limiting procedures to derive PS-PIDE models from on-lattice ABMs when the lattice parameters tend to zero (see e.g.  Macfarlane et al. 189, Lorenzi et al. 168). As a further note, structuring variables have also been incorporated into ‘mesoscale’ kinetic equations, which can also be scaled to a PDE form (see e.g. Erban and Othmer 88, Engwer et al. 85); we are, however, unaware of any such derivations that lead explicitly to the PS-PIDE form (2).
In the spirit of a tutorial-style example, here we formally derive a PS-PIDE model of form (2), starting from an ABM for the spatial spread and evolutionary dynamics of phenotype-structured cell populations (see Section 4.1.2). In this model, single cells undergo spatial movement, phenotypic changes, and proliferation and death according to a set of rules which correspond to a discrete-time branching random walk over a two-dimensional lattice [132], which represents one-dimensional physical and phenotypic domains (see Section 4.1.1). Note that here we will exemplify key ideas considering the case of unbounded domains, but both the governing rules for the dynamics of single cells in the ABM and the procedure employed to derive the corresponding PS-PIDE can easily be adapted to the case of bounded domains.
@@ 4.1 Tools and techniques to derive PS-PIDE models from agent-based models
Cell movement is modelled as a random walk along the spatial dimension. For a focal cell in the phenotypic state $y_{j}$ at spatial position $x_{i}$ at time $t_{k}$, the probability of moving left (i.e. to move to spatial position $x_{i-1}$) is $P^{L}_{x}(t_{k},x_{i},y_{j})$, the probability of moving right (i.e. to move to spatial position $x_{i+1}$) is $P^{R}_{x}(t_{k},x_{i},y_{j})$, and the probability of not moving (i.e. to remain stationary at position $x_{i}$) is then $P^{S}_{x}(t_{k},x_{i},y_{j})=1-P^{R}_{x}(t_{k},x_{i},y_{j})-P^{L}_{x}(t_{k},x_{i},y_{j})$. Here, $P^{L}_{x}$ and $P^{R}_{x}$ are non-negative real functions such that $P^{L}_{x}(t_{k},x_{i},y_{j})+P^{R}_{x}(t_{k},x_{i},y_{j})\leq 1$ for all $(t_{k},x_{i},y_{j})\in\mathbb{R}^{+}\times\mathbb{R}\times\mathbb{R}$.
@@
Phenotypic changes are incorporated into the model by allowing cells to update their phenotypic state according to a random walk along the phenotypic dimension. For a focal cell in the phenotypic state $y_{j}$ at spatial position $x_{i}$ at time $t_{k}$, the probability of entering the phenotypic state $y_{j-1}$ is $P^{L}_{y}(t_{k},x_{i},y_{j})$, the probability of entering the phenotypic state $y_{j+1}$ is $P^{R}_{y}(t_{k},x_{i},y_{j})$, and the probability of not undergoing phenotypic changes (i.e. remain in the phenotypic state $y_{j}$) is $P^{S}_{y}(t_{k},x_{i},y_{j})=1-P^{R}_{y}(t_{k},x_{i},y_{j})-P^{L}_{y}(t_{k},x_{i},y_{j})$. To reduce the amount of calculations in the derivation of the corresponding PS-PIDE model, we focus on the case where cells in the population undergo phenotypic changes with a constant probability, and thus we assume $P^{L}_{y}(t_{k},x_{i},y_{j})\equiv p^{L}_{y}\in\mathbb{R}^{+}$ and $P^{R}_{y}(t_{k},x_{i},y_{j})\equiv p^{R}_{y}\in\mathbb{R}^{+}$, which imply $P^{S}_{y}(t_{k},x_{i},y_{j})\equiv p^{S}_{y}=1-p^{R}_{y}-p^{L}_{y}$, with the parameters $p^{L}_{y}$ and $p^{R}_{y}$ such that $p^{L}_{y}+p^{R}_{y}\leq 1$.
@@
4.1.2 Formal derivation of the corresponding PS-PIDE model
@@
×$\;\times\Big{\{}p_{y}^{S}\,P^{S}_{x}(t,x,y)\,n(t,x,y)+$
$\;\;\;\,\,+p_{y}^{L}\,P^{S}_{x}(t,x,y+h_{y})\,n(t,x,y+h_{y})+p_{y}^{R}\,P^{S}_{x}(t,x,y-h_{y})\,n(t,x,y-h_{y})+$
$\;\;\;\,\,+p_{y}^{S}\,\Big{[}P^{R}_{x}(t,x-h_{x},y)\,n(t,x-h_{x},y)+P^{L}_{x}(t,x+h_{x},y)\,n(t,x+h_{x},y)\Big{]}+$
$\;\;\;\,\,+p_{y}^{L}\,\Big{[}P^{R}_{x}(t,x-h_{x},y+h_{y})\,n(t,x-h_{x},y+h_{y})+$
$+P^{L}_{x}(t,x+h_{x},y+h_{y})\,n(t,x+h_{x},y+h_{y})\Big{]}$
$\;\;\;\,\,+p_{y}^{R}\,\Big{[}P^{R}_{x}(t,x-h_{x},y-h_{y})\,n(t,x-h_{x},y-h_{y})+$
$+P^{L}_{x}(t,x+h_{x},y-h_{y})\,n(t,x+h_{x},y-h_{y})\Big{]}\Big{\}}\ .$
@@
Provided that the functions $P^{R}_{x}(t,x,y)$, $P^{L}_{x}(t,x,y)$, and $n(t,x,y)$ are sufficiently regular, one can substitute the first-order Taylor polynomial for $n(t+\tau,x,y)$ and the second-order Taylor polynomials for the other terms into the above equation. Dropping $(t,x,y)$ for brevity, after some simple (but tedious) calculations we find the following equation for the cell phenotype density $n$
@@
where higher order terms in $\tau$, $h_{x}$, and $h_{y}$ have been grouped into $h.o.t.\,$. If the functions $P^{L}_{x}$ and $P^{R}_{x}$ and the parameters $p^{L}_{y}$ and $p^{R}_{y}$ are also such that the following relations hold as $h_{x}\to 0$ and $h_{y}\to 0$
$P^{L}_{x}(t,x,y)-P^{R}_{x}(t,x,y)=\dfrac{h_{x}}{2}\Psi_{x}(t,x,y)+\mathcal{O}(h_{x}^{2})\ ,\quad\Psi_{x}:[0,\infty)\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}\ ,$
$P^{L}_{x}(t,x,y)+P^{R}_{x}(t,x,y)=\Phi_{x}(t,x,y)+\mathcal{O}(h_{x})\ ,\quad\Phi_{x}:[0,\infty)\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}^{+}\ ,$
@@
and recalling the definition of $\rho(t,x)$ given above, from the latter equation we find the following PS-PIDE model
@@
which is of the form of the PS-PIDE model (2), posed on a one-dimensional physical and phenotype domain, with
@@
Under different definitions of the functions $\Psi_{x}(t,x,y)$ and $\Phi_{x}(t,x,y)$, we can derive many of the models stated previously in Section 3, including:
@@
    from which, defining $D(y):=\alpha_{x}\hat{D}(y)$ and taking also $R(t,x,y,\rho)\equiv R(y,\rho)$, we obtain a PS-PIDE of form (7);
@@
    $\Psi_{x}(t,x,y):=-\,\hat{\mu}(y)\,\partial_{x}P(t,x),\qquad\Phi_{x}(t,x,y)\equiv 0\ ,\qquad P(t,x):=\Pi[\rho](t,x)\ ,$
    from which, defining $\mu(y):=\alpha_{x}\hat{\mu}(y)$ and taking $R(t,x,y,\rho)\equiv R(y,\Pi[\rho])\equiv R(y,P)$, we obtain a PS-PIDE of form (17);
@@
    $\Psi_{x}(t,x,y):=\hat{\chi}(y)\,\partial_{x}S(t,x)\ ,\qquad\Phi_{x}(t,x,y)\equiv\hat{D}\ ,$
    from which, defining $\chi(y):=\alpha_{x}\hat{\chi}(y)$ and $D:=\alpha_{x}\hat{D}$, and taking also $R(t,x,y,\rho)\equiv R(y,\rho,S(t,x))$, we obtain a PS-PIDE of form (20).
As a concluding remark we note that an excellent agreement can be found between the outputs of the ABM and the solutions of the corresponding PS-PIDE model, if the functions and parameters of the ABM are such that the conditions underlying the formal derivation carried out here are satisfied; an example is provided by the comparison between the plots in Figure 6(a) and Figure 6(b). This holds, in particular, when sufficiently large cell populations are considered. Then, demographic stochasticity (which cannot be captured by PS-PIDE models) does not play a dominant role in the ABM’s cell dynamics.
4.2 Tools and techniques to analyse PS-PIDE models
In the last decade, considerable attention has been directed towards analysing the behaviour of the solutions to PS-PIDE models of type (2), with particular emphasis on travelling waves and concentration phenomena. Referring back to Figures 6 and 7 – which display the results of numerical simulations of appropriately rescaled versions of the PS-PIDEs (17) and (20) – we can see phenotypic structuring across travelling waves as a result of the interplay between these two characteristic features: the cell densities $\rho(t,x)$ corresponding to solutions to these models behave like travelling waves (Figure 6(b), right panel and Figure 7, middle panels); within these waves, a concentration phenomenon occurs, with the phenotype density $n(t,x,y)$ being concentrated into a different phenotypic state $y$ at different spatial positions $x$ across the wave (Figure 6(b), left panel and Figure 7, left panels).
The obtained results, as well as the tools and techniques through which these results have been realised, build upon the mathematical literature on related PDE and non-local PDE models. Therefore, in order to make this section (to an extent) self-contained, before reaching PS-PIDE models (in Section 4.2.3), we first recall the essentials of travelling waves in PDE models of spatial spread dynamics (in Section 4.2.1) and concentration phenomena in non-local PDE models of evolutionary dynamics (in Section 4.2.2), taking the PDE (5) and the non-local PDE (10) as prototypical examples, respectively.

### v2

@@ 4 Mathematical tools and techniques
In this section, we provide a concise overview of mathematical tools and techniques that have been developed and deployed to: derive PS-PDE models of form (2) from agent-based models for the spatial spread and evolutionary dynamics of phenotype-structured cell populations (Section 4.1); analyse the qualitative properties of the solutions to such PS-PDE models (Section 4.2); and construct numerical solutions of these models (Section 4.3).
@@
4.1 Tools and techniques to derive PS-PDE models from agent-based models
Instead of defining a PS-PDE model of type (2) on the basis of population-scale phenomenological assumptions, it can be desirable to first postulate an ABM that tracks the spatial and evolutionary dynamics of individual cells, and then employ limiting procedures to derive the corresponding PS-PDE model [80]. In this way, the population-level terms that comprise the PS-PDE can be formally linked to explicit assumptions for the cellular processes that drive the dynamics at the single-cell level.
A range of approaches have been advanced and adopted in recent decades for transitioning between ABMs and PS-PDE models which describe simultaneously the spatial spread and evolutionary dynamics of phenotype-structured populations. These include probabilistic methods to derive PS-PDE models as the limit of corresponding off-lattice ABMs when the number of individuals in the population tends to infinity (see e.g.  Andrade-Restrepo et al 14, Champagnat and Méléard 52, Fontbona and Méléard 103, Leman 158), and formal limiting procedures to derive PS-PDE models from on-lattice ABMs when the lattice parameters tend to zero (see e.g.  Macfarlane et al 182, Lorenzi et al 173). As a further note, structuring variables have also been incorporated into ‘mesoscale’ kinetic equations, which can also be scaled to a PDE form (see e.g. Erban and Othmer 89, Engwer et al 87, Lorenzi et al 172); we are, however, unaware of any such derivations that lead explicitly to the PS-PDE form (2).
In the spirit of a tutorial-style example, here we formally derive a PS-PDE model of form (2), starting from an ABM for the spatial spread and evolutionary dynamics of phenotype-structured cell populations (see Section 4.1.2). In this model, single cells undergo spatial movement, phenotypic changes, and proliferation and death according to a set of rules which correspond to a discrete-time branching random walk over a two-dimensional lattice [135], which represents one-dimensional physical and phenotypic domains (see Section 4.1.1). Note that here we will exemplify key ideas considering the case of unbounded domains, but both the governing rules for the dynamics of single cells in the ABM and the procedure employed to derive the corresponding PS-PDE can easily be adapted to the case of bounded domains.
@@ 4.1 Tools and techniques to derive PS-PDE models from agent-based models
Cell movement is modelled as a random walk along the spatial dimension. For a focal cell in the phenotypic state $y_{j}$ at spatial position $x_{i}$ at time $t_{k}$, the probability of moving left (i.e. to spatial position $x_{i-1}$) is $P^{L}_{x}(t_{k},x_{i},y_{j})$, the probability of moving right (i.e. to spatial position $x_{i+1}$) is $P^{R}_{x}(t_{k},x_{i},y_{j})$, and the probability of not moving (i.e. remaining stationary at position $x_{i}$) is then $P^{S}_{x}(t_{k},x_{i},y_{j})=1-P^{R}_{x}(t_{k},x_{i},y_{j})-P^{L}_{x}(t_{k},x_{i},y_{j})$. Here, $P^{L}_{x}$ and $P^{R}_{x}$ are non-negative real functions such that $P^{L}_{x}(t_{k},x_{i},y_{j})+P^{R}_{x}(t_{k},x_{i},y_{j})\leq 1$ for all $(t_{k},x_{i},y_{j})\in\mathbb{R}^{+}\times\mathbb{R}\times\mathbb{R}$.
@@
Phenotypic changes are incorporated into the model by allowing cells to update their phenotypic state according to a random walk along the phenotypic dimension. For a focal cell in the phenotypic state $y_{j}$ at spatial position $x_{i}$ at time $t_{k}$, the probability of entering the phenotypic state $y_{j-1}$ is $P^{L}_{y}(t_{k},x_{i},y_{j})$, the probability of entering the phenotypic state $y_{j+1}$ is $P^{R}_{y}(t_{k},x_{i},y_{j})$, and the probability of not undergoing phenotypic changes (i.e. remaining in the phenotypic state $y_{j}$) is $P^{S}_{y}(t_{k},x_{i},y_{j})=1-P^{R}_{y}(t_{k},x_{i},y_{j})-P^{L}_{y}(t_{k},x_{i},y_{j})$. To reduce the amount of calculations in the derivation of the corresponding PS-PDE model, we focus on the case where cells in the population undergo phenotypic changes with a constant probability, and thus we assume $P^{L}_{y}(t_{k},x_{i},y_{j})\equiv p^{L}_{y}\in\mathbb{R}^{+}$ and $P^{R}_{y}(t_{k},x_{i},y_{j})\equiv p^{R}_{y}\in\mathbb{R}^{+}$, which imply $P^{S}_{y}(t_{k},x_{i},y_{j})\equiv p^{S}_{y}=1-p^{R}_{y}-p^{L}_{y}$, with the parameters $p^{L}_{y}$ and $p^{R}_{y}$ such that $p^{L}_{y}+p^{R}_{y}\leq 1$.
@@
4.1.2 Formal derivation of the corresponding PS-PDE model
@@
×$\;\times\Big{\{}\underbrace{p_{y}^{S}\,P^{S}_{x}(t,x,y)\,n(t,x,y)}_{\begin{subarray}{c}\mbox{\scriptsize{cells remaining at $(x,y)$}}\end{subarray}}+$
$\;\;\;\,\,+\underbrace{p_{y}^{L}\,P^{S}_{x}(t,x,y+h_{y})\,n(t,x,y+h_{y})+p_{y}^{R}\,P^{S}_{x}(t,x,y-h_{y})\,n(t,x,y-h_{y})}_{\begin{subarray}{c}\mbox{\scriptsize{cells switching from $(x,y\pm h_{y})$ to $(x,y)$}}\end{subarray}}+$
$\;\;\;\,\,+\underbrace{p_{y}^{S}\,\Big{[}P^{L}_{x}(t,x+h_{x},y)\,n(t,x+h_{x},y)+P^{R}_{x}(t,x-h_{x},y)\,n(t,x-h_{x},y)\Big{]}}_{\begin{subarray}{c}\mbox{\scriptsize{cells switching from $(x\pm h_{x},y)$ to $(x,y)$}}\end{subarray}}+$
$\;\;\;\,\,+\underbrace{p_{y}^{L}\,P^{R}_{x}(t,x-h_{x},y+h_{y})\,n(t,x-h_{x},y+h_{y})}_{\begin{subarray}{c}\mbox{\scriptsize{cells switching from $(x-h_{x},y+h_{y})$ to $(x,y)$}}\end{subarray}}+$
$\;\;\;\,\,+\underbrace{p_{y}^{L}\,P^{L}_{x}(t,x+h_{x},y+h_{y})\,n(t,x+h_{x},y+h_{y})}_{\begin{subarray}{c}\mbox{\scriptsize{cells switching from $(x+h_{x},y+h_{y})$ to $(x,y)$}}\end{subarray}}+$
$\;\;\;\,\,+\underbrace{p_{y}^{R}\,P^{R}_{x}(t,x-h_{x},y-h_{y})\,n(t,x-h_{x},y-h_{y})}_{\begin{subarray}{c}\mbox{\scriptsize{cells switching from $(x-h_{x},y-h_{y})$ to $(x,y)$}}\end{subarray}}+$
$\;\;\;\,\,+\underbrace{p_{y}^{R}\,P^{L}_{x}(t,x+h_{x},y-h_{y})\,n(t,x+h_{x},y-h_{y})}_{\begin{subarray}{c}\mbox{\scriptsize{cells switching from $(x+h_{x},y-h_{y})$ to $(x,y)$}}\end{subarray}}\Big{\}}\ .$
@@
Provided that the functions $P^{R}_{x}(t,x,y)$, $P^{L}_{x}(t,x,y)$, and $n(t,x,y)$ are sufficiently regular, one can substitute the first-order Taylor polynomial for $n(t+\tau,x,y)$ and the second-order Taylor polynomials for the other terms into the above equation. Dropping $(t,x,y)$ for brevity, after some simple (but tedious) calculations we are left with the following terms (arranged in increasing order of $\tau$, $h_{x}$, and $h_{y}$)
[…]
$\hskip 20.0pt+\dfrac{h_{x}^{2}}{2}\,\partial^{2}_{xx}\Big{[}\left(P^{R}_{x}+P^{L}_{x}\right)n\Big{]}+\dfrac{h_{y}^{2}}{2}\,\left(p^{R}_{y}+p^{L}_{y}\right)\,\partial^{2}_{yy}n+$
$\hskip 20.0pt+h_{x}h_{y}\,\left(p^{R}_{y}-p^{L}_{y}\right)\partial^{2}_{xy}\Big{[}\left(P^{R}_{x}-P^{L}_{x}\right)n\Big{]}+O(h_{x}^{2}h_{y})+O(h_{x}h_{y}^{2})\Big{\}}\,+$
$+\,\tau\,R(t,x,y,\rho)\times\Big{\{}\,n\,+O(h_{x})+O(h_{y})\,\Big{\}}+\ h.o.t.\,,$
[…]
@@
where higher order terms in $\tau$, $h_{x}$, and $h_{y}$ have again been grouped into $h.o.t.\,$. If the functions $P^{L}_{x}$ and $P^{R}_{x}$ and the parameters $p^{L}_{y}$ and $p^{R}_{y}$ are also such that the following relations hold as $h_{x}\to 0$ and $h_{y}\to 0$
$P^{L}_{x}(t,x,y)-P^{R}_{x}(t,x,y)=\dfrac{h_{x}}{2}\Psi_{x}(t,x,y)+\mathcal{O}(h_{x}^{2})\ ,\quad\Psi_{x}:\mathbb{R}^{+}\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}\ ,$
$P^{L}_{x}(t,x,y)+P^{R}_{x}(t,x,y)=\Phi_{x}(t,x,y)+\mathcal{O}(h_{x})\ ,\quad\Phi_{x}:\mathbb{R}^{+}\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}^{+}\ ,$
@@
and recalling the definition of $\rho(t,x)$ given above, from the latter equation we find the following PS-PDE model
@@
which is of the form of the PS-PDE model (2), posed on a one-dimensional unbounded physical and phenotype domain, with
@@
Choosing $\psi_{y}=0$, so that $\bar{A}=0$, under different definitions of the functions $\Psi_{x}(t,x,y)$ and $\Phi_{x}(t,x,y)$, we can derive many of the models stated previously in Section 3, including:
@@
    from which, defining $D(y):=\alpha_{x}\hat{D}(y)$ and taking also $R(t,x,y,\rho)\equiv R(y,\rho)$, we obtain a PS-PDE of form (7);
@@
    $\Psi_{x}(t,x,y):=\hat{\mu}(y)\,\partial_{x}P(t,x),\qquad\Phi_{x}(t,x,y)\equiv 0\ ,\qquad P(t,x):=\Pi[\rho](t,x)\ ,$
    from which, defining $\mu(y):=\alpha_{x}\hat{\mu}(y)$ and taking $R(t,x,y,\rho)\equiv R(y,\Pi[\rho])\equiv R(y,P)$, we obtain a PS-PDE of form (17);
@@
    $\Psi_{x}(t,x,y):=-\,\hat{\chi}(y)\,\partial_{x}S(t,x)\ ,\qquad\Phi_{x}(t,x,y)\equiv\hat{D}\ ,$
    from which, defining $\chi(y):=\alpha_{x}\hat{\chi}(y)$ and $D:=\alpha_{x}\hat{D}$, and taking also $R(t,x,y,\rho)\equiv R(y,\rho,S(t,x))$, we obtain a PS-PDE of form (20).
As a concluding remark we note that an excellent agreement can be found between the outputs of the ABM and the solutions of the corresponding PS-PDE model, if the functions and parameters of the ABM are such that the conditions underlying the formal derivation carried out here are satisfied; an example is provided by the comparison between the plots in Figure 6(a) and Figure 6(b). This holds, in particular, when sufficiently large cell populations are considered. Then, demographic stochasticity (which cannot be captured by PS-PDE models) does not play a dominant role in the ABM’s cell dynamics.
4.2 Tools and techniques to analyse PS-PDE models
In the last decade, considerable attention has been directed towards analysing the behaviour of the solutions to PS-PDE models of type (2), with particular emphasis on travelling waves and concentration phenomena. Referring back to Figures 6 and 7 – which display the results of numerical simulations of appropriately rescaled versions of the PS-PDEs (17) and (20) – we can see phenotypic structuring across travelling waves as a result of the interplay between these two characteristic features: the cell densities $\rho(t,x)$ corresponding to solutions to these models behave like travelling waves (Figure 6(b), right panel and Figure 7, middle panels); within these waves, a concentration phenomenon occurs, with the phenotype density $n(t,x,y)$ being concentrated into a different phenotypic state $y$ at different spatial positions $x$ across the wave (Figure 6(b), left panel and Figure 7, left panels).
The obtained results, as well as the tools and techniques through which these results have been realised, build upon the mathematical literature on related PDE and non-local PDE models. Therefore, in order to make this section (to an extent) self-contained, before reaching PS-PDE models (in Section 4.2.3), we first recall the essentials of travelling waves in PDE models of spatial spread dynamics (in Section 4.2.1) and concentration phenomena in non-local PDE models of evolutionary dynamics (in Section 4.2.2), taking the PDE (5) and the non-local PDE (10) as prototypical examples, respectively.